OpenStudy (anonymous):

Evaluate the function at the indicated values. f(x) = x^2 + 4x f(a)=?, f(-a)=?, f(1/a)=?

OpenStudy (anonymous):

simply substitute every time you see "x" in the function with the values a, -a, and 1/a. 1. f(a)=(a)^2+4(a) = a^2+4a 2. f(-a)=(-a)^2+4(-a) =a^2-4a 3. f(1/a)=(1/a)^2+4(1/a) = (1/a^2)+ (4/a)